Find 14.49 ÷ 7 Find 14.49 ÷ 7 Find 14.49 ÷ 7 Find 14.49 ÷ 7

Find 
 14.49 ÷ 7
Find 
 14.49 ÷ 7
Find 
 14.49 ÷ 7
Find 
 14.49 ÷ 7

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

We have,
= (1449/100) ÷ 7
Then,
= (1449/100) × (1/7)
= (207/100) × (1/1)
= (207/100)
On dividing a decimal by 100, the decimal point is shifted to the left by two places.
Then,
= 2.07

Thanks